OpenStudy (anonymous):

We have 100cm^2 of material to build a cylindrical can that will include the sides and bottom, but not the top. What are the dimensions of the can that will give the maximum enclosed volume?

OpenStudy (campbell_st):

so \[100 = 2\pi rh + \pi r^2\] then \[h = 50/(\pi r) - r/2\].........................(1) now volume of a cylinder is \[V = \pi r^2h\] so \[V = \pi r^2 (50/(\pi r) - r/2)\] or \[V = 50r - \pi r^3/2\] find dV/dr \[dV/dr = 50 - (3 \pi r^2)/2\] let dV/dr = 0 and solve for r \[r = \sqrt{100/3\pi}\] Substitute into (1) above to find h I think thats close to the solution

OpenStudy (anonymous):

I got it, thanks it actually solves quit nice in the end.